A particle moves along a straight line with an acceleration of a=4sinπt/2 m/s^2. If the displacement at t=0 is 0 and the initial velocity is -8/π m/s. find the acceleration after 2seconds and velocity after 2 seconds and displacement after 2seconds


1
Expert's answer
2021-08-30T15:05:08-0400

Given:

"a=4\sin\pi t/2"

"x_0=0"

"v_0=-8/\pi"


The acceleration

"a(2)=4\sin(\pi*2/2)=0\:\rm m/s^2"

The velosity

"v(t)=\int a(t)dt=\int 4\sin\pi t/2dt \\=-8/\pi\cos\pi t/2+C,\quad C=0"

"v(2)=-8/\pi\cos(\pi *2/2)=8/\pi"

The displacement

"x(t)=\int v(t)dt=\int (-8/\pi\cos\pi t/2)dt"

"=-16/\pi\sin\pi t/2"

"x(2)=-16/\pi\sin(\pi*2/2)=0"